# Optimal Control of Reserve Asset Portfolios for Pegged Digital Currencies

### v1

@@ Abstract
Stablecoins promise par convertibility, yet issuers must balance immediate liquidity against yield on reserves to keep the peg credible. We study this treasury problem as a continuous-time control task with two instruments: reallocating reserves between cash and short-duration government bills, and setting a spread fee for either minting or burning the coin. Mint and redemption flows follow mutually exciting processes that reproduce clustered order flow; peg deviations arise when redemptions exceed liquid reserves within settlement windows. We develop a stochastic model predictive control framework that incorporates moment closure for event intensities. Using Pontryagin’s Maximum Principle, we demonstrate that the optimal control exhibits a bang-off-bang structure: each asset type is purchased at maximum capacity when the utility difference exceeds the corresponding difference in shadow costs. Introducing settlement windows leads to a sampled-data implementation with a simple threshold (soft-thresholding) structure for rebalancing. We also establish a monotone stress-response property: as expected outflows intensify or windows lengthen, the optimal policy shifts predictably toward cash. In simulations covering various stress test scenarios, the controller preserves most bill carry in calm markets, builds cash quickly when stress emerges, and avoids unnecessary rotations under transitory signals. The proposed policy is implementation-ready and aligns naturally with operational cut-offs. Our results translate empirical flow risk into auditable treasury rules that improve peg quality without sacrificing avoidable carry.
@@
Stablecoin; peg stability; reserve management; stochastic optimal control; model predictive control; hawkes processes; liquidity risk; cryptocurrency
@@ 1 Introduction
While fiat-backed stablecoins operate outside the full decentralization paradigm, their treasury management problem intersects with broader dynamics in decentralized finance (DeFi). A literature review by [2] identifies stablecoins as a central class of ”financial tokens” in DeFi, emphasizing that even fully collateralized designs can deviate from their peg, exhibit state-dependent volatility, and depend critically on liquidity provision and external data feeds. This positioning situates issuer reserve policies alongside on-chain mechanisms, market microstructure effects, and inter-protocol linkages that can amplify or dampen peg pressures.
Related market-design research links validator and block-builder incentives to liquidity and price stability in on-chain asset markets. [3] show that block-production timing strategies can yield measurable gains in maximal extractable value (MEV) without necessarily harming consensus stability. For pegged digital assets that rely on on-chain liquidity pools, MEV-maximizing behaviors among market participants can affect execution quality for redemptions and rebalancing, introducing another stochastic element into to the issuer’s reserve-management problem.
@@ 2.2.1 Demand And Peg Stability
$\lambda_{R}(t)$ $=\lambda_{0,R}+\int_{0}^{t}\kappa_{R}e^{-\theta_{R}(t-s)}\mathrm{d}N_{R}(s)+\zeta\Delta P(t),$ (1)
$\lambda_{M}(t)$ $=\lambda_{0,M}+\int_{0}^{t}\kappa_{M}e^{-\theta_{M}(t-s)}\mathrm{d}N_{M}(s),$ (2)
where $\lambda_{0,R},\lambda_{0,M}\geq 0$ are baseline intensities, $\kappa_{R},\kappa_{M}>0$ capture self-excitation (each event increases the likelihood of subsequent events), and $\theta_{R},\theta_{M}>0$ govern the exponential decay of this excitation. The redemption intensity includes an additional feedback term $\zeta\Delta P(t)^{2}$ with $\zeta\geq 0$, reflecting how peg instability accelerates redemptions. This mechanism endogenously models bank-run dynamics characterized by self-reinforcing distrust in the currency’s stability.
@@
$\mathrm{d}R(t)=\int_{\mathbb{R}_{+}}zN_{R}(\mathrm{d}t,\mathrm{d}z),\quad\mathrm{d}M(t)=\int_{\mathbb{R}_{+}}zN_{M}(\mathrm{d}t,\mathrm{d}z),$ (3)
where marks $z$ represent individual transaction sizes with finite first moments.
The peg deviation evolves deterministically according to unmet redemption demand relative to outstanding supply:
$\frac{\mathrm{d}}{\mathrm{d}t}\Delta P(t)\;=\;\eta\frac{[Q(t)-R_{\mathrm{liq}}(t)]_{+}}{S_{\mathrm{out}}(t)}\;-\;\gamma\delta(t),$ (4)
where $R_{\mathrm{liq}}(t)$ denotes immediately available cash reserves, $S_{\mathrm{out}}(t)=S_{\mathrm{out}}(0)+M(t)-R(t)$ is the outstanding supply, and $\delta(t)$ represents the mint/burn fee spread. The deviations resulting from equation (4) directly reflect deficits in reserve coverage, rather than being mitigated by supply and demand. The first term captures how liquidity shortfalls—when redemption requests $Q(t)$ exceed available cash $R_{\mathrm{liq}}(t)$—create peg pressure proportional to the fraction of supply that cannot be serviced. The parameter $\eta>0$ governs this sensitivity, while $\gamma>0$ determines how effectively fees can counteract peg deviations. This deterministic relationship makes peg stability an operational challenge: maintaining $\Delta P(t)\approx 0$ requires either sufficient liquidity buffers or appropriate fee policies to offset redemption pressure.
@@ 2.2.2 Reserve-Balance State Model
The state constraints $R_{\mathrm{liq}}(t)\geq 0$ and $R_{\mathrm{bill}}(t)\geq 0$, combined with the bounded reallocation rate $|\omega(t)|\leq\omega_{\max}$, endogenize the fundamental liquidity-yield trade-off: higher Treasury allocations increase returns but reduce the buffer against redemption surges, potentially triggering peg instability through the mechanism in equation (4). The balance sheet identity $R_{\mathrm{liq}}(t)+R_{\mathrm{bill}}(t)=S_{\mathrm{out}}(t)$ (up to small operational items) ensures conservation of reserves relative to outstanding supply.
@@ 2.2.3 Optimal Control Problem
$\dot{\Delta P}(t)$ $=\eta\frac{[Q(t)-R_{\mathrm{liq}}(t)]_{+}}{S_{\mathrm{out}}(t)}-\gamma\delta(t),$ (13)
@@
Admissible controls must satisfy the operational constraints (6), be progressively measurable with respect to the filtration $\{\mathcal{F}_{t}\}$, and maintain state feasibility $R_{\mathrm{liq}}(t),R_{\mathrm{bill}}(t)\geq 0$.
@@
The infinite-horizon problem faces well-posedness challenges from cost integral convergence despite potentially unbounded Treasury growth, the non-smooth term $Q(t)-R_{\mathrm{liq}}(t)]_{+}$ in peg dynamics, and potentially explosive paths from the peg-redemption feedback loop. The following standard conditions ensure well-posedness: (i) positive discount rate $\rho>0$ with bounded interest rates; (ii) subcritical Hawkes processes satisfying $\kappa_{R}/\theta_{R}<1$ and $\kappa_{M}/\theta_{M}<1$ ensuring stationary intensities; (iii) bounded controls maintaining nonnegative reserves; and (iv) the balance sheet constraint preventing unbounded carry extraction.
@@ 3.1 Stochastic Model Predictive Control
where the self-excitation terms $\kappa_{X}\hat{\lambda}_{X}(t)$ arise from taking expectations of future jump contributions. These ODEs exploit the Markovian property of exponential Hawkes processes: given current intensities, the expected future evolution depends only on these values, not the full history. The expected net redemption flow $\hat{Q}(t)=\bar{z}_{R}\hat{\lambda}_{R}(t)-\bar{z}_{M}\hat{\lambda}_{M}(t)$ then enters the deterministic system dynamics, with the crucial feedback that peg deviations influence future redemption intensity through the $\zeta\Delta P(t)^{2}$ term. This coupling necessitates joint propagation of all state variables, as flows and peg trajectory depend on each other dynamically. Given mean transaction sizes $\bar{z}_{R}$ and $\bar{z}_{M}$, the expected net redemption flow becomes $\hat{Q}(t)=\bar{z}_{R}\hat{\lambda}_{R}(t)-\bar{z}_{M}\hat{\lambda}_{M}(t)$. Substituting these deterministic forecasts into the system dynamics yields
@@
$\dot{\Delta P}(t)$ $=\eta\frac{[\hat{Q}(t)-R_{\mathrm{liq}}(t)]_{+}}{\hat{S}_{\mathrm{out}}(t)}-\gamma\delta(t),$ (23)
@@
3.2 Event-Triggered Optimization via Pontryagin’s Maximum Principle

### v2

# Optimal Control of Reserve Asset Portfolios for Stablecoins

@@ Abstract
Stablecoins promise par convertibility, yet issuers must balance immediate liquidity against yield on reserves to keep the peg credible. We study this treasury problem as a continuous-time control task with two instruments: reallocating reserves between cash and short-duration government bills, and setting a spread fee for either minting or burning the coin. Mint and redemption flows follow mutually exciting processes that reproduce clustered order flow. Peg deviations arise when immediate cash coverage is insufficient relative to outstanding supply, and the market price relaxes toward this liquidity-coverage fair value. We develop a stochastic model predictive control framework that incorporates moment closure for event intensities. Using Pontryagin’s Maximum Principle, we show that the optimal reallocation control exhibits a soft-thresholding structure: no rebalancing occurs when the shadow-cost differential lies within a deadzone set by transaction costs, and reallocation scales linearly beyond that threshold up to a capacity-imposed saturation limit. Introducing settlement windows leads to a sampled-data implementation with a simple threshold (soft-thresholding) structure for rebalancing. We also establish a monotone stress-response property: as expected outflows intensify or windows lengthen, the optimal policy shifts predictably toward cash. In simulations covering various stress test scenarios, the controller preserves most bill carry in calm markets, builds cash quickly when stress emerges, and avoids unnecessary rotations under transitory signals. The proposed policy is implementation-ready and aligns naturally with operational cut-offs. Our results translate empirical flow risk into auditable treasury rules that improve peg quality without sacrificing avoidable carry.
@@
Stablecoin; peg stability; reserve management; stochastic optimal control; model predictive control; Hawkes processes
@@ 1 Introduction
Closest to the present paper is [8], who develop a discrete-time dynamic model of a fiat-backed stablecoin issuer that allocates reserves between cash and bonds to maximize discounted cash flows. They show that liquidity and capital requirements jointly reduce default probability and the expected price impact of forced bond sales, and provide a calibrated mapping from prudential targets to requirement levels. Our work differs in that we adopt the issuer’s operational perspective rather than solving for a regulatory optimum, and in the modelling assumptions on both market dynamics and the control framework. While the framework of [8] informs the calibration of prudential requirements, ours translates such requirements into auditable, real-time treasury rules. While fiat-backed stablecoins operate outside the full decentralization paradigm, their treasury management problem intersects with broader dynamics in decentralized finance (DeFi). A literature review by [18] identifies stablecoins as a central class of ”financial tokens” in DeFi, emphasizing that even fully collateralized designs can deviate from their peg, exhibit state-dependent volatility, and depend critically on liquidity provision and external data feeds. This positioning situates issuer reserve policies alongside on-chain mechanisms, market microstructure effects, and inter-protocol linkages that can amplify or dampen peg pressures.
Related market-design research links validator and block-builder incentives to liquidity and price stability in on-chain asset markets. [21] show that block-production timing strategies can yield measurable gains in maximal extractable value (MEV) without necessarily harming consensus stability. For pegged digital assets that rely on on-chain liquidity pools, MEV-maximizing behaviors among market participants can affect execution quality for redemptions and rebalancing, introducing another stochastic element to the issuer’s reserve-management problem.
@@
An analysis in [6] documents substantial heterogeneity in reserve practices among major stablecoin issuers in 2021: Tether managed roughly USD 63 billion across a diversified set of instruments with heterogeneous liquidity, whereas USDC held about USD 32 billion entirely in cash; Diem, by contrast, recommended holding  10% in cash and  90% in short-term government securities. This work extends both the academic literature and current treasury practice by developing a control-theoretic reserve-management framework that adapts the portfolio in real time to observed flows and yields while respecting current legislation and asset-quality constraints. Our policy replaces static, rule-based heuristics with an implementable, settlement-aware, closed-loop design that (i) delivers transparent, window-based rebalancing rules with explicit execution frictions and (ii) maximizes expected carry subject to quantified de-peg penalties and hard cash-availability constraints.
@@ 2.1 Regulatory Requirements
Throughout, we model cash as operationally available on demand. In real treasury operations, ”cash” often means balances at commercial banks or custodians and can therefore carry counterparty and access risk. This is consistent with episodes where a stablecoin’s secondary-market price deviated from par due to perceived custodial risk even when the issuer remained solvent. Operationally, issuers mitigate this by (i) diversifying transaction-account balances across multiple depository institutions, (ii) preferring central-bank accounts where available, (iii) sweeping excess cash into Treasury-only vehicles with same-day liquidity, and (iv) maintaining committed liquidity lines for intraday settlement shocks.
@@ 2.2.1 Demand And Peg Stability
$\lambda_{R}(t)$ $=\lambda_{0,R}+\int_{0}^{t}\kappa_{R}e^{-\theta_{R}(t-s)}\,\mathrm{d}N_{R}(s)+\int_{0}^{t}\zeta\,e^{-\theta_{R}(t-s)}\Delta P(s)^{2}\,\mathrm{d}s,$ (1)
$\lambda_{M}(t)$ $=\lambda_{0,M}+\int_{0}^{t}\kappa_{M}e^{-\theta_{M}(t-s)}\,\mathrm{d}N_{M}(s).$ (2)
where $\lambda_{0,R},\lambda_{0,M}\geq 0$ are baseline intensities, $\kappa_{R},\kappa_{M}>0$ capture self-excitation (each event increases the likelihood of subsequent events), and $\theta_{R},\theta_{M}>0$ govern the exponential decay of this excitation. The redemption intensity includes an additional feedback term driven by $\Delta P(s)^{2}$, reflecting how peg instability accelerates redemptions. This mechanism endogenously models bank-run dynamics characterized by self-reinforcing distrust in the currency’s stability.
@@
$\mathrm{d}M(t)=\int_{\mathbb{R}_{+}}z\,N_{M}(\mathrm{d}t,\mathrm{d}z),\qquad\mathrm{d}R(t)=\int_{\mathbb{R}_{+}}\min\bigl(z,\,S_{\mathrm{out}}(t^{-})\bigr)\,N_{R}(\mathrm{d}t,\mathrm{d}z),$ (3)
where marks $z$ represent individual transaction sizes with finite first moments. To rule out mechanically infeasible paths in which cumulative executed redemptions exceed the outstanding supply, we truncate each redemption mark at the remaining supply. The outstanding supply evolves as
$S_{\mathrm{out}}(t)=S_{\mathrm{out}}(0)+M(t)-R(t),\qquad S_{\mathrm{out}}(0)>0,$ (4)
where the left-limit $S_{\mathrm{out}}(t^{-})$ in (3) ensures that executed redemptions satisfy $R(t)\leq S_{\mathrm{out}}(0)+M(t)$ and hence $S_{\mathrm{out}}(t)\geq 0$ for all $t$.
The conditional expected net cash outflow rate is
$\hat{Q}(t):=\mathbb{E}\!\left[Q(t)\mid\mathcal{F}_{t}\right]=\bar{z}_{R}\,\lambda_{R}(t)-\bar{z}_{M}\,\lambda_{M}(t),$ (5)
where $\bar{z}_{R}$ and $\bar{z}_{M}$ denote the mean redemption and mint sizes, respectively. Over a settlement horizon $\tau_{\mathrm{set}}>0$, the quantity $\tau_{\mathrm{set}}\,[\hat{Q}(t)]_{+}$ represents the expected cumulative cash drain that remains after netting mints against redemptions. Comparing this drain with the immediately available cash reserve $R_{\mathrm{liq}}(t)$ yields a normalized liquidity shortfall
$\Delta P_{\mathrm{fair}}(t):=\mathbf{1}_{\{S_{\mathrm{out}}(t)>0\}}\left[\frac{\tau_{\mathrm{set}}\,[\hat{Q}(t)]_{+}-R_{\mathrm{liq}}(t)}{S_{\mathrm{out}}(t)}\right]_{0}^{1},$ (6)
where $x]_{+}:=\max\{x,0\}$ and $x]_{0}^{1}:=\min\{1,\max\{x,0\}\}$. The quantity $\Delta P_{\mathrm{fair}}(t)$ measures the fraction of outstanding supply whose redemption cannot be settled from current liquid reserves within the relevant horizon. We interpret $\Delta P_{\mathrm{fair}}(t)$ as the fair-value peg deviation that would prevail if the secondary-market price instantaneously reflected this settlement-liquidity gap.
The actual peg deviation then evolves deterministically through first-order mean reversion toward $\Delta P_{\mathrm{fair}}(t)$ at rate $\eta>0$, supplemented by a corrective term proportional to the mint/burn fee spread $\delta(t)$:
$\dot{\Delta P}(t)=\eta\!\Big(\Delta P_{\mathrm{fair}}(t)-\Delta P(t)\Big)-\gamma\,\delta(t).$ (7)
The parameter $\eta>0$ governs this sensitivity, while $\gamma>0$ determines how effectively fees can counteract peg deviations. This deterministic relationship makes peg stability an operational challenge: maintaining $\Delta P(t)\approx 0$ requires either sufficient liquidity buffers or appropriate fee policies to offset redemption pressure.
@@ 2.2.2 Reserve-Balance State Model
We impose the constraint $R_{\mathrm{bill}}(t)\geq 0$ for all $t$, since the bill bucket cannot be negative. In contrast, the liquid cash position can be temporarily negative in practice due to intraday settlement mismatches and/or the use of committed liquidity lines. Accordingly, we allow $R_{\mathrm{liq}}(t)\in\mathbb{R}$. Combined with the bounded reallocation rate $|\omega(t)|\leq\omega_{\max}$, the constraints endogenize the fundamental liquidity-yield trade-off: higher Treasury allocations increase returns but reduce the buffer against redemption surges, potentially triggering peg instability through the mechanism in equation (7). The reserve base satisfies $R_{\mathrm{liq}}(t)+R_{\mathrm{bill}}(t)\geq S_{\mathrm{out}}(t)$ for all $t\geq 0$ whenever the inequality holds at $t=0$, since interest accrual on Treasury holdings generates an excess $r_{\mathrm{bill}}\,R_{\mathrm{bill}}\geq 0$, which means the issuer is overcollateralized by the cumulative interest earned.
@@ 2.2.3 Optimal Control Problem
[…]
@@
Admissible controls must satisfy the operational constraints (9), be progressively measurable with respect to the filtration $\{\mathcal{F}_{t}\}$, and maintain state feasibility with $R_{\mathrm{bill}}(t)\geq 0$.
@@
The infinite-horizon optimization problem faces several well-posedness challenges: convergence of the cost integral despite potentially unbounded Treasury growth, the non-smooth clipping in the fair-value mapping $\Delta P_{\mathrm{fair}}(t)$ defined in (6) appearing in the peg dynamics, and potential explosive paths arising from the peg-redemption feedback loop. We address these challenges through the following conditions: (i) a positive discount rate $\rho>0$ with bounded interest rates; (ii) subcritical Hawkes processes satisfying $\kappa_{R}/\theta_{R}<1$ and $\kappa_{M}/\theta_{M}<1$, which guarantee stationary intensities and finite first moments; (iii) bounded controls that maintain nonnegative reserves; and (iv) the balance sheet constraint that prevents unbounded carry extraction by tying reserves to outstanding supply.
A crucial observation is that the problem admits a finite-value admissible policy, which serves as a witness for well-posedness. Since every outstanding coin is backed one-to-one, we can construct a peg-neutral policy with finite discounted cost. Consider the all-cash policy defined by
$u^{\mathrm{cash}}(t)=(\omega(t),\delta(t))\equiv(0,0),\quad R_{\mathrm{bill}}(t)\equiv 0,\quad R_{\mathrm{liq}}(t)=S_{\mathrm{out}}(t),$ (19)
where the issuer holds the full reserve in immediately available cash without trading between buckets or charging fees.
Under this all-cash policy, immediately available cash equals outstanding supply at all times, so
$\Delta P_{\mathrm{fair}}(t)=\left[1-\frac{R_{\mathrm{liq}}(t)}{S_{\mathrm{out}}(t)}\right]_{+}=\left[1-\frac{S_{\mathrm{out}}(t)}{S_{\mathrm{out}}(t)}\right]_{+}=0.$
The peg dynamics therefore reduce to
$\dot{\Delta P}(t)=-\eta\,\Delta P(t)-\gamma\,\delta(t).$
With $\delta(t)\equiv 0$ and the natural initialization $\Delta P(0)=0$, we obtain $\Delta P(t)\equiv 0$ for all $t\geq 0$.
For this admissible trajectory, the running cost simplifies to
$\ell\bigl(x(t),u^{\mathrm{cash}}(t)\bigr)=c_{\mathrm{peg}}\cdot 0+c_{\mathrm{fee}}\cdot 0+\lambda_{\omega}\cdot 0+\tfrac{1}{2}\rho_{\omega}\cdot 0-r_{\mathrm{bill}}(t)\cdot 0=0,$ (20)
and consequently the discounted infinite-horizon objective equals
$J_{\infty}\bigl(x(0);u^{\mathrm{cash}}\bigr)=\int_{0}^{\infty}e^{-\rho t}\cdot 0\,dt=0<\infty.$ (21)
Therefore, the value function is finite at $x(0)$ and the optimization problem is proper: the infimum of the discounted cost is not $\infty$ and a minimizing sequence exists. This witness policy demonstrates why the peg-redemption feedback cannot cause explosive costs in our formulation - there exists an admissible control that nullifies the fair-value deviation term $\Delta P_{\mathrm{fair}}$ in the peg law, keeping $\Delta P$ identically zero with no associated cost.
Together with the existence of the admissible peg-neutral policy demonstrated above, these conditions ensure that the discounted infinite-horizon problem is well-posed: the value function remains finite for the initial state, and the state-control constraints admit at least one trajectory with bounded cost. This establishes an economically meaningful baseline of perfect convertibility with zero peg error, against which all other policies can be evaluated within the same discounted framework.
2.2.4 Assumptions and robustness discussion
The model is deliberately stylized to yield an auditable, settlement-aware feedback rule rather than a fully structural description of stablecoin secondary-market pricing. This subsection discusses which assumptions are primarily adopted for tractability, how they can affect quantitative outputs, and how they may be relaxed while preserving the qualitative policy logic (threshold rebalancing and monotone stress response):
  1. 1.
    Flow model (Hawkes) and calibration. We use exponential Hawkes processes because they reproduce clustered order flow with a Markovian intensity state, which is what enables the moment-closure forecast and the MPC implementation. Empirically, the Hawkes parameters $(\lambda_{0,\cdot},\kappa_{\cdot},\theta_{\cdot})$ can in principle be estimated directly from on-chain mint/burn timestamps and updated over rolling windows. These estimates can be fed into the filtered intensity updates (23)-(24). We emphasize that the degree of self-excitation (branching ratios $\kappa/\theta$) is quantitatively important because it affects the speed and persistence of clustered redemptions and therefore the optimal liquidity buffer. Our scenario design partially addresses this sensitivity by varying both baseline intensities and moving the redemption branching ratio close to criticality, which should be interpreted as a robustness check over plausible flow regimes rather than a point-calibrated empirical claim.
  2. 2.
    Objective specification and sensitivity. The stage cost (12) is intended as a reduced-form aggregation of (i) peg-quality and reputational costs, (ii) user-friction costs from fees, (iii) trading frictions/market impact, and (iv) bill carry. The calibration in Section 4.2 should be interpreted as illustrative scaling chosen to match operational tolerances. While alternative penalty choices would shift numerical thresholds and cash targets, the qualitative control form is robust to these variations. The framework is designed to be re-calibrated to each issuer’s mandate rather than used with universal weights.
  3. 3.
    Interpretation and limitation (confidence-driven de-pegs). Equation (7) models an operational liquidity-driven component of de-pegging: $\Delta P$ increases when redemptions exceed immediately available settlement liquidity within the relevant window. Historical episodes suggest that secondary-market prices can also deviate from par due to anticipatory fear even when redemptions remain available. Our formulation intentionally isolates the liquidity/settlement channel because it is the part most directly controllable by treasury allocation and execution timing.
@@ 3.1 Stochastic Model Predictive Control
where the self-excitation terms $\kappa_{X}\hat{\lambda}_{X}(t)$ arise from taking expectations of future jump contributions. These ODEs exploit the Markovian property of exponential Hawkes processes: given current intensities, the expected future evolution depends only on these values, not the full history. The expected outstanding supply within the MPC horizon evolves as
$\dot{\hat{S}}_{\mathrm{out}}(t)=\bar{z}_{M}\hat{\lambda}_{M}(t)-\bar{z}_{R}\hat{\lambda}_{R}(t)=-\hat{Q}(t),\qquad\hat{S}_{\mathrm{out}}(t_{k})=S_{\mathrm{out}}(t_{k}).$ (27)
Substituting these deterministic forecasts into the system dynamics yields
@@
$\dot{\Delta P}(t)$ $=\eta\big(\hat{\Delta P}_{\mathrm{fair}}(t)-\Delta P(t)\big)-\gamma\,\delta(t),$ (30)
$\hat{\Delta P}_{\mathrm{fair}}(t)$ $:=\mathbf{1}_{\{\hat{S}_{\mathrm{out}}(t)>0\}}\left[\frac{\tau_{\mathrm{set}}\,[\hat{Q}(t)]_{+}-R_{\mathrm{liq}}(t)}{\hat{S}_{\mathrm{out}}(t)}\right]_{0}^{1}.$ (31)
@@
3.2 Optimal Control Characterization via Pontryagin’s Maximum Principle